OpenStudy (anonymous):

system of elimination 2x-2y=8 x+2y=-1

OpenStudy (bahrom7893):

add the two equations. What do you get?

OpenStudy (anonymous):

these are simultaneous equations. You need to solve for one variable first then substitute it into the second equation in the ordered set to solve for the other equation and you will end up with an (x,y) point value.

OpenStudy (anonymous):

In this case take the first equation and set y=0 2x-2y =8 2x-2(0) =8 2x =8 x =8/2 x =4 Substitute this into the second equation... x+2y=-1 4+2y=-1 2y=-5 y=-5/2 therefore this ordered set has the answer of:  Hope that helps :)